OpenStudy (anonymous):

f(x)=4 times 3^x+5-6

OpenStudy (anonymous):

\[f(x)=4 \times 3^{x+5}-6\]

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

domain?

OpenStudy (anonymous):

what exactly do you have to do?

OpenStudy (anonymous):

domain range and asymptote

OpenStudy (anonymous):

domain (-infinity,infinity)

myininaya (myininaya):

the inverse of y=ln(x) is y=e^x which looks like this on a graph (y=ln(x) is a reflection of y=e^x about y=x)  oops i don't have to explain the domain you got the domain great so the range now so the range for y=e^x is (0,inf) right like we can't get a negative output same thing for y=a^x where a>0 and not equal to 1 y=a^x has domain (-inf,inf) and range (0,inf) now y=(a^x)+1 has domain (-inf,inf) and range (1,inf) since each point on y=a^x was raised up once...